November 15th 2024

From Hall A Wiki
Jump to: navigation, search
  • Updates
    • CDet
      • retroffiting of all modules : 1.5 left to add straps and remount - by end of November
      • HV cables setup , need 2 HV cards
      • first set of 168 installed, 2nd setup almost done - plan to be done by end of December
      • will look at LV today and move in hall - Test power all NINO with new supply, if ok with supplu will build distribution
      • DAQ crate : all TDC are in -
      • might not be able to take cosmics before mounting detector
      • will try to take cosmics with module before depending on cabling
    • HCal
      • connected trigger cables to TI
      • waiting to setup stand alone
    • GEMs
      • could take cosmics data during week-end : trigger cable was cut
      • block size was changed : 4 MB - will check evio block size
      • GEMs were uncabled - to move SBS -
      • Back tracker testing need all VTPs crates ~ 3 weeks
      • histogramming method : need to be tested and maybe optimized , test in INFN test - check Chandan
      • MPD transceiver : IC2 avalaible - need to modify driver and VTP firmware to handle that
    • ECal
      • cosmics trigger ready : to be deployed
      • cluster still need to be done : December
      • all VXS crates are in
      • missing VTP and SD
      • will connect network next week for ECal crates and VTP
    • RF time
      • 1190 need to be installed for copy
    • Beamline : need to check when Paul is coming